# bibliography/references.bib *[itemize]labelindent=10pt, itemindent=0pt, leftmargin=* TrackFormers: In Search of Transformer-Based Particle Tracking for the High-Luminosity LHC Era

### v2

@@ 1.1. Our contribution
Our goal is to offer fast, i.e., computationally-efficient and high-throughput, solutions for tracking. Therefore, we strive for models that assign hits to trajectories in a single inference step. We determine the accuracy and speed of these models for the tracking problem and use simulated data of increasing complexity to compare our approaches. Finally, we discuss how our initial approaches can improve their performance through post-processing steps and argue that our approaches are fast and accurate enough to be considered as candidates for future tracking pipelines. Our data sets (Odyurt:2024:DATASET) and code (Dobreva:2024:CODE)111Also available at https://github.com/TrackFormers/TrackFormers-Models are publicly available.
@@ 1.2.5. Advances in deep learning and GNNs
Given that particle interactions can be conveniently represented as graphs, Graph Neural Networks (GNNs) have emerged as a promising approach. These networks operate on graph-structured data, capturing the relational information between hits in a detector. The state-of-the-art in recent years has focused on GNN-based solutions (Bakina:2022:DLTR; Biscarat:2021:TRTR; Choma:2020:TSLE; DeZoort:2021:CPTE; Duarte:2022:GNNP; Elabd:2022:GNNC; Farrell:2018:NDLM; Heintz:2020:ACPT; Ju:2020:GNNP; Ju:2021:PGDL; Mieskolainen:2023:HNCH; Murnane:2023:EGNN; Thais:2021:ISGO; Duarte:2022:GNNP; Caillou:2024:PPAG), or see this reference for a review on GNNs in HEP (Shlomi:2021:GNNP).
@@ 2.1. REDVID data set
For more detailed descriptions covering field data types, we refer readers to the README file in the shared Zenodo record (Odyurt:2024:DATASET)
@@ 2.2. TrackML data set
The TrackML data set is a simulated collection of proton-proton collision events. It simulates specifically the production of top quark-antiquark pairs in the high-pileup conditions expected at the HL-LHC, with each event incorporating an average of 200 pileup interactions. The simulated detector is designed as a generalized representation of an LHC tracker, featuring discrete layers of cylinder and disk-shaped sensor arrays operating within a magnetic field as layers. The data set provides 3D hit positions along with truth-level information about the particles responsible for generating these hits. Some simplifications with respect to more realistic data were made. For example, it does not account for reconstructed hit merging, as such occurrences are rare (less than 0.5%). Moreover, it excludes complex smaller detector elements such as electronic systems, cooling tubes, or cables.
@@
Though resulting from a simulation, the TrackML data has considerable scale. It contains 8 850 individual events with an average of 10 780 particles, between 6 898 and 16 197, per event, which could be interpreted as the average track count. It includes particles represented with as many as 28 hit points. At the same time, there are particles with 0 associated hit points. The average particle hit count throughout the events is 8.52. The number of individual particle vertices composing a bunch crossing varies from 1300 to 3048. Substantial amount of noise hits are present (between 18 590 and 20 518 per event), which is one of the main factors making TrackML a challenging data set. Almost every publication considering this data set has performed some sort of noise reduction. For our highest complexity data set, “200-500 tracks per event from the TrackML data set”, we have selected between 200 to 500 tracks at random per event. This process is repeated five times, resulting in 43 725 reduced events.
The TrackML data set is described extensively in literature, for further information please refer to (Kiehn:2019:TrackML) or (Amrouche:2020:TMLC).
@@ 3.1. Transformer designs
Typically, attention is computed on a set of queries at the same time, making , and matrices. Multi-Head Attention (MHA) allows for jointly attending to information from different representation sub-spaces at different positions in the data. MHA is summerised by
for which is the number of heads, is a parameter matrix, with and as the dimensionality of the embedding. The attention heads are used in parallel with dimensionality – this optimisation reduces the computational cost, making it comparable to that of single-head attention. Overall, the shape of the matrices is then , with batch size , number of attention heads , and sequence length .
@@ 3.2. U-Net design
The U-Net (Ronneberger:2015:UNet) is a CNN architecture that primarily segments images. It aims to classify each pixel in an image and allows it to distinguish and separate objects in complicated visual data. The ”U” in U-Net represents its U-shaped structure: it has a contracting path (encoder) and a wide-ranging path (decoder) with a bottleneck in between. On the one hand, contextual information is caught by the contracting path through the progressive downsampling of the input image using convolutional layers and pooling operations that reduce the spatial dimensions and increase feature abstraction. This process allows the network to grasp the global structure of objects in the image while improving its analytical capabilities. The most abstracted features are extracted and the smallest resolution is reached at the bottleneck. On the other hand, the decoder part aims to reconstruct the image at the original resolution by employing transposed convolutions or upsampling layers. It gradually refines the output by combining contextual information from the encoder with finer details provided by the skip connections thereby ensuring that the spatial details lost during downsampling are satisfactorily recovered.
@@ 3.3. Design choices
The input sequence to the Transformers is a full event, with the tokens corresponding to hit coordinates (x,y,z), which we do not discretise. As a result of the unordered nature of the recorded hits, we do not use positional encoding for the Transformers. For models involving an encoder-only design, i.e., EncCla and EncReg, padding is used to allow variable length inputs. Note that different events could, for instance, have variable track counts, thus resulting in variable numbers of hits. The following sections will elaborate each model design and approach in detail. A simple depiction is presented in Figure 2, covering Transformer-based model approaches, i.e., EncDec (Figure 2(a)), EncCla (Figure 2(b)) and EncReg (Figure 2(c)).
@@ 3.3.4. Model 4 - U-Net
The last model implemented has supposed an alternative methodology to Transformer-like models. Due to the nature of the pixel regression task under consideration, the utilisation of a U-Net-based model appears to be a suitable option for a fruitful methodology. Vanilla U-Net models usually consider a set of the ordinary well-known dense convolutional layers, which are the core block building the network and where the major part of the computational effort occurs. For this use case, the input data should be preprocessed into a multi-dimensional tensor of size , i.e. a three-dimensional tensor encompassing only one channel which contains discrete values of 0 and 1, indicating whether there is background or hit. Due to the sparse nature of this information, a more suitable U-Net model can be build if the convolutional operations are considered to occur directly on the sparse domain, thereby ignoring the background data governing the overall tensor representing the events. Consequently, since the hit occupancy for a certain event will be very low in general, the convolutional layers have been substituted by their sparse option (Graham:2017:SSCN; Graham:2018:3DSS), thereby giving name to the sparse U-Net model that has been considered.
@@
Perhaps the most challenging aspect of this U-Net-based approach is the considered preprocessing. In the preprocessing step, when converting the original set of numerical values to the set of to create the sparse tensor of coordinates, a scale factor different from one can be considered. In addition, also during this step, it is possible to carry out an interpolation procedure to add more hits between the original ones. The scaling factor determines the physical spacing between the original points in the three-dimensional data. Conversely, the number of points used for interpolation between consecutive points enhances the amount of information provided to the network. Notably, these parameters were considered exclusively during the training process, leaving the test data unaltered. Different interpolation methods can be employed. In this case, the cubic interpolation method from the SciPy library has been utilised. As a consequence, both factors will introduce hyperparameters that need exploration, since a bigger scale factor will demand a higher amount of interpolated data, and vice-versa. However, a fixed scale factor of and also interpolated hits between each pair of original coordinates have lead to the best results that are presented in this manuscript.
@@
Regarding the hyperparameters of the entire pipeline, one may expect the performance to vary with respect to different aspects. The first one refers to a certain arbitrary factor of scale that can be considered during the process of converting from the set to the one, while the second one would represent the considered amount of interpolated hits for each event. The largest model trained has a total of about 8 000 000 parameters.
@@ 4.1. Prediction performance
To measure the prediction accuracy of our model designs, we consider a custom metric, named FitAccuracy score. It is essentially identical to the TrackML score (Kiehn:2019:TrackML), with a small modification that concerns the case of REDVID-generated data sets only. The TrackML score is calculated based on the association of weights to hits in the TrackML data. However, REDVID simulations do not generate particles and thus do not consider weights, but tracks. Essentially, we replace the true particle with the true track and consider the weight value of 1 for all hits to arrive at our custom FitAccuracy value. In this fashion, we can have a single comparable scoring for all data sets and model designs. In the definition of the TrackML score and, consequently, FitAccuracy, reconstructed tracks with four or more hits are considered, and at least 50% of a reconstructed track’s hits must originate from the same truth particle for that track to be considered for the scoring. The score of a track is the sum of correctly assigned hit weights. Available scoring for each model is provided in Table 1.
@@ 4.2. Computational performance
Table 2 can be used to infer a rough estimate for the scaling of the mean inference GPU-time as a function of the input size. However, a more detailed analysis was conducted for the EncReg-FA model, trained on the 200–500 tracks data set. This model is particularly well-suited for scaling to larger events due to its integration of FlashAttention. The results, presented in Figure 5, demonstrate that while the inference GPU-time does not scale linearly with input size, its growth remains sub-quadratic. Notably, even for events with 8 000 tracks, the mean inference time remains below 200 ms, which is promising for applications to high-luminosity LHC data, where similar track multiplicities are anticipated. It is worth noting, however, that for events with higher track counts, larger models might offer improved performance, albeit at the cost of increased execution time. Consequently, this study should be viewed as a rough demonstration of scaling under the specific constraint of fixed model size.
@@ 5. Discussion - Applying the method
The U-Net model can also determine the track classes for all hits in one step. The inference time increases slightly with larger input matrices and model sizes. However, it was not possible to achieve good accuracy for simulations with higher complexity levels (above level 3). This method was proposed as an alternative to Transformer-based models with attention mechanisms, but its performance was found to depend on parameters such as input image scale and the number of interpolation points in the training set. The REDVID data sets did not require sparse convolutional networks, but this approach ensured scalability for TrackML data. In addition, implementation restrictions of attention mechanisms within thespconv (Spconv:2022:SSCL) package hindered direct comparisons with Transformers. This limitation restricts the UNet model’s applicability to less complex scenarios.
@@ 6. Conclusion and future work
Another conclusion is the effectiveness of our method where we consider a complexity spectrum for the problem ranging from a simple to a complex reconstruction problem. Both the prediction accuracy and the computational effort results from Figure 6 provide metrics for the early elimination of underperforming machine learning designs. It goes without saying that the design and validation effort (both human and machine) for lower complexity levels is incomparably lower than when tackling the problem in its full complexity.

### v3

@@ 1.1. Our contribution
Our goal is to offer fast, i.e., computationally-efficient and high-throughput, solutions for tracking. Therefore, we strive for models that assign hits to trajectories in a single inference step. We determine the accuracy and speed of these models for the tracking problem and use simulated data of increasing complexity to compare our approaches. Finally, we discuss how our initial approaches can improve their performance through post-processing steps and argue that our approaches are fast and accurate enough to be considered as candidates for future tracking pipelines. Our data sets (Odyurt:2024:DATASET) and code (Dobreva:2024:CODE)111Also available at: https://github.com/TrackFormers/TrackFormers-Models are publicly available.
@@ 1.2.5. Advances in deep learning and GNNs
Given that particle interactions can be conveniently represented as graphs, Graph Neural Networks (GNNs) have emerged as a promising approach. These networks operate on graph-structured data, capturing the relational information between hits in a detector. The state-of-the-art in recent years has focused on GNN-based solutions (Bakina:2022:DLTR; Biscarat:2021:TRTR; Choma:2020:TSLE; DeZoort:2021:CPTE; Duarte:2022:GNNP; Elabd:2022:GNNC; Farrell:2018:NDLM; Heintz:2020:ACPT; Ju:2020:GNNP; Ju:2021:PGDL; Mieskolainen:2023:HNCH; Murnane:2023:EGNN; Thais:2021:ISGO; Caillou:2024:PPAG), or see this reference for a review on GNNs in HEP (Shlomi:2021:GNNP).
@@ 2.1. REDVID data set
For more detailed descriptions covering field data types, we refer readers to the README file in the shared Zenodo record (Odyurt:2024:DATASET).
@@ 2.2. TrackML data set
The TrackML data set is a simulated collection of proton-proton collision events. It simulates specifically the production of top quark-antiquark pairs in the high-pileup conditions expected at the HL-LHC, with each event incorporating an average of 200 pileup interactions. The simulated detector is designed as a generalised representation of an LHC tracker, featuring discrete layers of cylinder and disk-shaped sensor arrays operating within a magnetic field as layers. The data set provides 3D hit positions along with truth-level information about the particles responsible for generating these hits. Some simplifications with respect to more realistic data were made. For example, it does not account for reconstructed hit merging, as such occurrences are rare (less than 0.5%). Moreover, it excludes complex smaller detector elements such as electronic systems, cooling tubes, or cables.
@@
Though resulting from a simulation, the TrackML data has considerable scale. It contains 8 850 individual events with an average of 10 780 particles, between 6 898 and 16 197, per event, which could be interpreted as the average track count. It includes particles represented with as many as 28 hit points. At the same time, there are particles with 0 associated hit points. The average particle hit count throughout the events is 8.52. The number of individual particle vertices composing a bunch crossing varies from 1300 to 3048. Substantial amount of noise hits are present (between 18 590 and 20 518 per event), which is one of the main factors making TrackML a challenging data set. Almost every publication considering this data set has performed some sort of noise reduction. For our highest complexity data set, “200-500 tracks per event from the TrackML data set”, we have selected between 200 to 500 tracks at random per event. This process is repeated five times, resulting in 43 725 reduced events. The tracks and hits are selected completely at random without any special attention to, e.g., number of noisy hits, i.e., we do not apply any noise reduction.
The TrackML data set is described extensively in literature. For further information please refer to (Kiehn:2019:TrackML) or (Amrouche:2020:TMLC).
@@ 3.1. Transformer designs
Typically, attention is computed on a set of queries at the same time, making , and matrices. Multi-Head Attention (MHA) allows for jointly attending to information from different representation sub-spaces at different positions in the data. MHA is summarised by
for which is the number of heads, is a parameter matrix, with and as the dimensionality of the embedding. The attention heads are used in parallel with dimensionality – this optimisation reduces the computational cost, making it comparable to that of single-head attention. Overall, the shape of the matrices is then , equivalent to , with batch size , number of attention heads , and sequence length .
@@ 3.2. U-Net design
The U-Net (Ronneberger:2015:UNet) is a CNN architecture that primarily segments images. It aims to classify each pixel in an image and allows it to distinguish and separate objects in complicated visual data. The “U” in U-Net represents its U-shaped structure: it has a contracting path (encoder) and a wide-ranging path (decoder) with a bottleneck in between. On the one hand, contextual information is caught by the contracting path through the progressive downsampling of the input image using convolutional layers and pooling operations that reduce the spatial dimensions and increase feature abstraction. This process allows the network to grasp the global structure of objects in the image while improving its analytical capabilities. The most abstracted features are extracted and the smallest resolution is reached at the bottleneck. On the other hand, the decoder part aims to reconstruct the image at the original resolution by employing transposed convolutions or upsampling layers. It gradually refines the output by combining contextual information from the encoder with finer details provided by the skip connections thereby ensuring that the spatial details lost during downsampling are satisfactorily recovered.
@@ 3.3. Design choices
The input sequence to the Transformers is a full event, with the tokens corresponding to hit coordinates , which we do not discretise. As a result of the unordered nature of the recorded hits, we do not use positional encoding for the Transformers. For models involving an encoder-only design, i.e., EncCla and EncReg, padding is used to allow variable length inputs. Note that different events could, for instance, have variable track counts, thus resulting in variable numbers of hits. The following sections will elaborate each model design and approach in detail. A simple depiction is presented in Figure 2, covering Transformer-based model approaches, i.e., EncDec (Figure 2(a)), EncCla (Figure 2(b)) and EncReg (Figure 2(c)).
@@ 3.3.4. Model 4 - U-Net
The last model implemented has supposed an alternative methodology to Transformer-like models. Due to the nature of the pixel regression task under consideration, the utilisation of a U-Net-based model appears to be a suitable option for a fruitful methodology. Vanilla U-Net models usually consider a set of the ordinary well-known dense convolutional layers, which are the core block building the network and where the major part of the computational effort occurs. For this use case, the input data should be preprocessed into a multi-dimensional tensor of size , i.e., a three-dimensional tensor encompassing only one channel which contains discrete values of 0 and 1, indicating whether there is background or hit. Due to the sparse nature of this information, a more suitable U-Net model can be build if the convolutional operations are considered to occur directly on the sparse domain, thereby ignoring the background data governing the overall tensor representing the events. Consequently, since the hit occupancy for a certain event will be very low in general, the convolutional layers have been substituted by their sparse option (Graham:2017:SSCN; Graham:2018:3DSS), thereby giving name to the sparse U-Net model that has been considered.
@@
Perhaps the most challenging aspect of this U-Net-based approach is the considered preprocessing. In the preprocessing step, when converting the original set of numerical values to the set of to create the sparse tensor of coordinates, a scale factor different from one can be considered. In addition, also during this step, it is possible to carry out an interpolation procedure to add more hits between the original ones. The scaling factor determines the physical spacing between the original points in the three-dimensional data. Conversely, the number of points used for interpolation between consecutive points enhances the amount of information provided to the network. Notably, these parameters were considered exclusively during the training process, leaving the test data unaltered. Different interpolation methods can be employed. In this case, the cubic interpolation method from the SciPy library has been utilised. As a consequence, both factors will introduce hyperparameters that need exploration, since a bigger scale factor will demand a higher amount of interpolated data, and vice-versa. However, a fixed scale factor of and also interpolated hits between each pair of original coordinates have led to the best results that are presented in this manuscript.
@@
Regarding the hyperparameters of the entire pipeline, one may expect the performance to vary with respect to different aspects. The first one refers to a certain arbitrary factor of scale that can be considered during the process of converting from the set to the one, while the second one would represent the considered amount of interpolated hits for each event. The largest model trained has a total of about 8 000 000 parameters. The said hyperparameters influence data processing but not the model itself. Since a balanced approach is desired, the training data was formatted as three-dimensional tensors for U-Net processing, with flexibility in scaling and interpolation. Various scaling factors (, , , ) and interpolation levels (5, 10, 20, 50 points) were tested. A scale with 10 interpolated points per hit provided the best trade-off between performance and training efficiency. However, it is important to note that these hyperparameters are not a priority, as they pertain to data preprocessing rather than the neural network itself.
@@ 4.1. Prediction performance
To measure the prediction accuracy of our model designs, we consider a custom metric, named FitAccuracy score. It is essentially identical to the TrackML score (Kiehn:2019:TrackML), with a small modification that concerns the case of REDVID-generated data sets only. The TrackML score is calculated based on the association of weights to hits in the TrackML data. However, REDVID simulations do not generate particles and thus do not consider weights, but tracks. Essentially, we replace the true particle with the true track and consider the weight value of 1 for all hits to arrive at our custom FitAccuracy value. In this fashion, we can have a single comparable scoring for all data sets and model designs. In the definition of the TrackML score and, consequently, FitAccuracy, reconstructed tracks with four or more hits are considered, and at least 50% of a reconstructed track’s hits must originate from the same truth particle for that track to be considered for the scoring. The score of a track is the sum of correctly assigned hit weights. As such, the FitAcurracy score gives an indication of the fraction of correctly reconstructed tracks, and can thus be interpreted as a (weighted) reconstruction efficiency. Available scoring for each model is provided in Table 1.
@@ 4.2. Computational performance
Table 2 can be used to infer a rough estimate for the scaling of the mean inference GPU-time as a function of the input size. However, a more detailed analysis was conducted for the EncReg-FA model, trained on the 200–500 tracks data set. This model is particularly well-suited for scaling to larger events due to its integration of FlashAttention. The results, presented in Figure 5, demonstrate that while the inference GPU-time does not scale linearly with input size, its growth remains sub-quadratic. Notably, even for events with 8 000 tracks, the mean inference time remains below 200 ms, which is promising for applications to high-luminosity LHC data, where similar track multiplicities are anticipated. It is worth noting, however, that for events with higher track counts, larger models might offer improved physics performance, albeit at the cost of increased execution time. Consequently, this study should be viewed as a rough demonstration of scaling under the specific constraint of fixed model size. We do recognise that input scale is only one dimension of complexity and Figure 5 intentionally does not take into account model prediction qualities, i.e., physics accuracy. The main point conveyed is that the indicated increase in inference computational burden is superior to traditional algorithms.
@@ 5. Discussion - Applying the method
The U-Net model can also determine the track classes for all hits in one step. The inference time increases slightly with larger input matrices and model sizes. However, it was not possible to achieve good accuracy for simulations with higher complexity levels (above level 3). This method was proposed as an alternative to Transformer-based models with attention mechanisms, but its performance was found to depend on parameters such as input image scale and the number of interpolation points in the training set. The REDVID data sets did not require sparse convolutional networks, but this approach ensured scalability for TrackML data. In addition, implementation restrictions of attention mechanisms within the spconv (Spconv:2022:SSCL) package hindered direct comparisons with Transformers. This limitation restricts the U-Net model’s applicability to less complex scenarios.
@@
While negotiating the full scale and complexity of a realistic detector data set requires further research, our methodology provides an effective approach. To recapitulate, instead of picking one ML model design and iteratively improving it, we have opted for multiple designs. The layered approach to complexity allows for speedy evaluation and elimination of bad designs, making the design journey rather cost-effective. The way forward from this point on is to move forward in the complexity spectrum, represented by data sets of higher complexity and higher scale. In this particular case, models EncCla and EncReg, or improved variations of these shall be considered. Further fine-tuning along the way, in the form of post-processing models or computational steps could be considered. In short, we have the opportunity to try many design candidates, precisely because we can do it in a cost-effective way.
@@ 6. Conclusion and future work
Another conclusion is the effectiveness of our method where we consider a complexity spectrum for the problem ranging from a simple to a complex reconstruction problem. Both the prediction accuracy and the computational effort results from Figure 6 provide metrics for the early elimination of underperforming machine learning designs. It goes without saying that the design and validation effort (both human and machine) for lower complexity levels is incomparably lower than when tackling the problem in its full complexity. It should be noted also that the most complex data set considered in this work is still an order of magnitude removed from the true HL-LHC data in hit multiplicity. The scaling behaviour with increasingly complex data sets, as shown in Figure 6 and Table 2, looks promising in particular for the EncCla model and to a lesser extent, for the EncReg model. However, a full evaluation of the included models up to the complexity of HL-LHC data is considered follow-up work, as significant adaptations would be required including the integration of pre- and post-processing steps into the algorithm, which may substantially optimise performance. Furthermore, computational developments, such as parallelised training and inference of the Transformer models may also enable more efficient scaling.